A population consists of values (1, 4, 7). Consider all possible samples of size n= 3 that can be drawn without replacement from this population.

a. Find the mean of the population.

b. Find the standard deviation of the population.

c. Find the mean of the sampling distribution of means.

d. Find the standard deviation of the sample distribution of means,

e.Construct the probability histogram of x with replacement


1
Expert's answer
2022-04-27T14:39:40-0400

a. We have population values 1,4,7, population size N=3 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{1+4+7}{3}=4"

b. Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{3}(9+0+9)=6"


"\sigma=\sqrt{\sigma^2}=\sqrt{6}\approx2.4495"

c. The number of possible samples which can be drawn with replacement is "N^n=3^3=27"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 &1,1,1 & 1\\
 \hdashline
 2 & 1,1,4 & 2 \\
 \hdashline
 3 & 1,1,7 & 3\\
 \hdashline
 4 & 1,4,1 & 2\\
 \hdashline
 5 & 1,4,4 & 3 \\
 \hdashline
 6 & 1,4,7& 4 \\
 \hdashline
 7 & 1,7,1 & 3 \\
 \hdashline
 8 & 1,7,4 & 4 \\
 \hdashline
 9 & 1,7,7 & 5 \\
 \hdashline
 10 & 4,1,1 & 2\\
 \hdashline
 11 & 4,1,4 & 3 \\
 \hdashline 
 12 & 4,1,7 & 4 \\
 \hdashline 
13 & 4,4,1 & 3 \\
 \hdashline 
 14 & 4,4,4 & 4 \\
 \hdashline 
 15 & 4,4,7 & 5 \\
 \hdashline 
 16 & 4,7,1 & 4 \\
 \hdashline 
 17 & 4,7,4 & 5 \\
 \hdashline 
 18 & 4,7,7 & 6\\
 \hdashline 
 19 & 7,1,1 & 3\\
 \hdashline
 20 & 7,1,4 & 4\\
 \hdashline 
 21 & 7,1,7 & 5\\
 \hdashline 
 22 & 7,4,1 & 4\\
 \hdashline
 23 & 7,4,4 & 5\\
 \hdashline 
 24 & 7,4,7 & 6\\
 \hdashline
 25 & 7,7,1 & 5\\
 \hdashline 
 26 & 7,7,4 & 6\\
 \hdashline 
 27 & 7,7,7 & 7\\
 \hdashline 
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 1 & 1/27 & 1/27 & 1/27 \\
 \hdashline
 2 & 3/27 & 6/27 & 12/27 \\
 \hdashline
 3 & 6/27 & 18/27 & 54/27 \\
 \hdashline
 4 & 7/27 & 28/27 & 112/27 \\
 \hdashline
 5 & 6/27 & 30/27 & 150/27 \\
 \hdashline
 6 & 3/27 & 18/27 & 108/27 \\
 \hdashline
 7 & 1/27 & 7/27 & 49/27 \\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=4=\mu"



d. The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{486}{27}-(4)^2=2= \dfrac{\sigma^2}{n}"

"\sigma_{\bar{X}}=\sqrt{2}\approx1.4142"

e.